What steps are needed to graph the lines lines: y = 0.15x +0.79 : y = 0.11x - y = -0.85

Answer by stanbon(75887)

What steps are needed to graph the lines :
y = 0.15x +0.79
You only need two points:
Let x = 0, then y = 0.79
Let x = 1, the y = 0.79+0.15 = 0.94
---------------------------------------
y = 0.11x - y = -0.85
Too many equal signs.
======================================
Cheers,
Stan H.